OpenStudy (anonymous):

What are the solutions of x2 + 2x + 9 = 0?

OpenStudy (anonymous):

There are no soluions

OpenStudy (anonymous):

the solution is a complex no.

OpenStudy (anonymous):

When you try the quadratic formula, you will have a negative number under the square root part: \[\frac{-b \pm \sqrt{b^2-4ac}}{2a}=\frac{-2\pm\sqrt{4-36}}{2}=\frac{-2\pm\sqrt{-32}}{2}\] (This is equivalent to saying the discriminant is negative). As a result, there are no real solutions to this equation.

OpenStudy (anonymous):

Solve x2 – 3x = 1

OpenStudy (anonymous):

bring the 1 to the other side, and use the same procedure as above: \[x^2-3x=1 \ \ \ \Longleftrightarrow \ \ \ x^2-3x-1=0\]